SubtoNOT:

What is the simplified product of z_(1) and z_(2) given that z_(1)=8+3i and z_(2)=1-4i? differences too

hannahj:

U dont deserve help

hannahj:

but here 20-29i

SubtoNOT:

And differences (i really do)

hannahj:

no more help for you bud

SubtoNOT:

HELP NOW MEAN

Jayisbetter:

To find the product of complex numbers � 1 = 8 + 3 � z 1 ​ =8+3i and � 2 = 1 − 4 � z 2 ​ =1−4i, you can use the distributive property of multiplication. Let's calculate the product � 1 ⋅ � 2 z 1 ​ ⋅z 2 ​ : � 1 ⋅ � 2 = ( 8 + 3 � ) ⋅ ( 1 − 4 � ) = 8 ⋅ 1 + 8 ⋅ ( − 4 � ) + 3 � ⋅ 1 + 3 � ⋅ ( − 4 � ) = 8 − 32 � + 3 � − 12 � 2 z 1 ​ ⋅z 2 ​ ​ =(8+3i)⋅(1−4i) =8⋅1+8⋅(−4i)+3i⋅1+3i⋅(−4i) =8−32i+3i−12i 2 ​ Remember that � 2 = − 1 i 2 =−1, so you can substitute that in: 8 − 32 � + 3 � − 12 � 2 = 8 − 32 � + 3 � + 12 = ( 8 + 12 ) + ( − 32 � + 3 � ) = 20 − 29 � 8−32i+3i−12i 2 ​ =8−32i+3i+12 =(8+12)+(−32i+3i) =20−29i ​ So, the product � 1 ⋅ � 2 z 1 ​ ⋅z 2 ​ is 20 − 29 � 20−29i. The real part is 20, and the imaginary part is -29.

Jayisbetter:

oopd

SubtoNOT:

id find

hannahj:

@jayisbetter wrote:
To find the product of complex numbers � 1 = 8 + 3 � z 1 ​ =8+3i and � 2 = 1 − 4 � z 2 ​ =1−4i, you can use the distributive property of multiplication. Let’s calculate the product � 1 ⋅ � 2 z 1 ​ ⋅z 2 ​ : � 1 ⋅ � 2 = ( 8 + 3 � ) ⋅ ( 1 − 4 � ) = 8 ⋅ 1 + 8 ⋅ ( − 4 � ) + 3 � ⋅ 1 + 3 � ⋅ ( − 4 � ) = 8 − 32 � + 3 � − 12 � 2 z 1 ​ ⋅z 2 ​ ​ =(8+3i)⋅(1−4i) =8⋅1+8⋅(−4i)+3i⋅1+3i⋅(−4i) =8−32i+3i−12i 2 ​ Remember that � 2 = − 1 i 2 =−1, so you can substitute that in: 8 − 32 � + 3 � − 12 � 2 = 8 − 32 � + 3 � + 12 = ( 8 + 12 ) + ( − 32 � + 3 � ) = 20 − 29 � 8−32i+3i−12i 2 ​ =8−32i+3i+12 =(8+12)+(−32i+3i) =20−29i ​ So, the product � 1 ⋅ � 2 z 1 ​ ⋅z 2 ​ is 20 − 29 � 20−29i. The real part is 20, and the imaginary part is -29.
lmao

SubtoNOT:

???????????????????????

SubtoNOT:

i solved it alreadyyyyy =(

hannahj:

@subtonot wrote:
i solved it alreadyyyyy =(
Good job

Jayisbetter:

real

YourlocalRandom2000:

z_1 \times z_2 = (8 + 3i) \times (1 - 4i) ] Using the distributive property and the fact that ( i^2 = -1 ), we can expand and simplify this expression: [ z_1 \times z_2 = 8 \times 1 - 8 \times 4i + 3i \times 1 - 3i \times 4i ] [ z_1 \times z_2 = 8 - 32i + 3i - 12i^2 ] [ z_1 \times z_2 = 8 - 32i + 3i + 12 ] [ z_1 \times z_2 = 20 - 29i ] So, the simplified product of ( z_1 ) and ( z_2 ) is ( 20 - 29i ).

hannahj:

@yourlocalrandom2000 wrote:
z_1 \times z_2 = (8 + 3i) \times (1 - 4i) ] Using the distributive property and the fact that ( i^2 = -1 ), we can expand and simplify this expression: [ z_1 \times z_2 = 8 \times 1 - 8 \times 4i + 3i \times 1 - 3i \times 4i ] [ z_1 \times z_2 = 8 - 32i + 3i - 12i^2 ] [ z_1 \times z_2 = 8 - 32i + 3i + 12 ] [ z_1 \times z_2 = 20 - 29i ] So, the simplified product of ( z_1 ) and ( z_2 ) is ( 20 - 29i ).
Post elsewhere

Testacc:

amgnsla

Testacc:

1 attachment
SubtoNOT:

?????

@testacc wrote: